A high-performance sports car can go from 0 to 100 mph in 5.6 s. (Assume the car travels in the positive direction. Indicate the direction with the sign of your answer.)

(a) What is the car's average acceleration (in SI units; meter-kilogram-second)?

(b) The same car can come to a complete stop from 31 m/s in 5.9 s. What is its average acceleration?
Answer #1

a) Avg accn = (44.704-0)/5.6 = 7.983 m/s^2

b) Avg accn. = (0-31)/5.9 = -5.254 m/s^2
Answer #2

100 mph = 44.704 m/s


a) accelration = 44.704/ 5.6 = 7.98 m/s^2


b) average acceleration = 0-31 / 5.9 =5.25 m/s^2
Answer #3

(a)the car's average acceleration is

a = (v - u/t)

where v = 100 mph = (100/3600) m/s and t = 5.6 s

b)the initial and final speeds of the car are u = 31 m/s and v = 0 in time t = 5.9 s

the average acceleration is

a = (v - u/t)
Answer #4

(a) average acceleration = (100 - 0) / 5.6 * (4/9) m/s^2

= 4000/504 = 7.9365 m/s^2

(b) average acceleration = (0 - 31) / 5.9 m/s^2

= -310/59 = -5.254 m/s^2

the negative sign shows that the acceleration is in opposite direction or its deaccelerating.